# Implementation of time-dependent Hartree Fock in real space

### v1

@@ Abstract
Time-dependent Hartree-Fock (TDHF) is one of the fundamental post-Hartree-Fock (HF) methods to describe excited states. In its Tamm-Dancoff form, equivalent to Configuration Interaction Singles, it is still widely used and particularly applicable to big molecules where more accurate methods may be unfeasibly expensive. However, it is rarely implemented in real space, mostly because of the expensive nature of the exact-exchange potential in real space. Compared to widely used Gaussian-type orbitals (GTO) basis sets, real space often offers easier implementation of equations and more systematic convergence of Rydberg states, as well as favorable scaling, effective domain parallelization, flexible boundary conditions, and ability to treat model systems. We implemented TDHF in the Octopus real-space code as a step toward linear-response hybrid time-dependent density-functional theory (TDDFT), other post-HF methods, and ensemble density-functional theory methods involving exact exchange. Calculation of HF’s non-local exact exchange is very expensive in real space. We overcome this limitation with Octopus’ implementation of Adaptively Compressed Exchange (ACE), and find the appropriate mixing and starting point to complete the ground-state calculation in a practical amount of time, to enable TDHF. We compared our results to those from GTOs on a set of small molecules and confirmed close agreement of results, though with larger deviations than in the case of semi-local TDDFT. We find that convergence of TDHF demands a finer real-space grid than semi-local TDDFT. We also present the subtleties in benchmarking a real-space calculation against GTOs, relating to Rydberg and vacuum states.
@@ 1 Introduction
Hartree-Fock (HF) theory [1, 2] is a main pillar in quantum chemistry, serving as the starting point for calculating the electronic structure of molecules. HF calculations provide molecular geometry, electron distribution, and ground-state properties [3]. Once the ground state has been calculated, excited states can be calculated using the ground-state results. A first approximation one can make for the excitation energies is from the eigenvalue differences of the unoccupied and occupied states, termed the Independent Particle Approximation (IPA) [4]. Configuration Interaction (CI), Equation-of-Motion Coupled Cluster (EOM-CC), and Time-Dependent Hartree-Fock (TDHF) are examples of post-Hartree-Fock methods [5] that can provide more accurate excited states. The exact exchange potential used in HF is orbital-dependent while it is density-dependent in traditional semi-local Kohn-Sham Density Functional Theory (DFT). Hybrid functionals which are a mixture of these two types of exchange potentials are widely used [6, 7, 8] to improve accuracy, so TDHF is a building block for time-dependent density-functional theory (TDDFT) with hybrids.
@@
Hartree-Fock theory provides a reasonably accurate description of the ground state of molecular systems, especially for small-to-moderately sized molecules and systems where electron correlation effects are not dominant [9]. Gaussian-type orbitals (GTOs) are typically used as basis sets for Hartree-Fock [10]. Implementation of Hartree-Fock in real space is rare, though a finite-element-type implementation has been published [11]. Using GTOs conveniently allows integrals to be done analytically. However, the nature of orbitals one can obtain in an electronic structure calculations heavily depends on the type and size of the basis set used. The basis set is typically optimized for the ground state which can make description of unoccupied orbitals difficult [12] This issue is particularly acute for Rydberg states which are highly excited states that are far from the atoms and weakly bound [13], but transitions from valence to Rydberg states can be important in optical properties [14]. To include these states, additional diffuse functions is added to the basis sets [15]. Comparisons of GTOs and real space have shown that such very large, diffuse basis sets can be needed to describe the excited states involved in non-linear optical response [16]. In real space, Rydberg states do not require special diffuse functions but can be captured by systematically converging the physical size of the real-space domain in which the orbitals are calculated. the number of grid-points which can be systematically converged [17]. Other advantages of real-space calculations include favorable scaling for large systems, effective parallelization over domains, and straightforward use of both finite and periodic boundary conditions [17, 18, 19], and ability to handle non-atomic model systems [20]. In real-space however, the calculation of exchange integrals becomes very expensive [21] and these integrals are a major part of any HF calculation underlying TDHF.
@@
In this paper, we present our results from our real-space implementation of TDHF in the Octopus code. In section 2, we discuss the general theory behind ground-state HF and TDHF. Section 3.1 discusses the exact exchange potential which is the main feature of HF which sets it apart from traditional Kohn-Sham DFT. As this is the most expensive part of the calculation, we discuss a formalism, ACE, which has been available in Octopus for a while now for hybrid DFT calculations to speed up the exact exchange calculation. We also present the general parameters used in our calculations and the need for them in sections 3.3 and 3.4. In section 4 we discuss the general difficulties in reaching a solution to the HF ground-state in real-space and how to overcome them. Finally, section 4.5 shows the comparison of excitation energies obtained from GTO-based approach to the real-space approach, demonstrating a good level of agreement and practical attainability of convergence with respect to the real-space grid parameters.
@@ 2 Theory
The four terms in the bracket are kinetic energy operator, external potential, Hartree potential, and exchange potential. The Hartree potential depends on the density $\rho\left(r\right)=\sum_{j}^{N_{e}}\left|\psi_{j}\right|^{2}$, where $N_{e}$ is the number of electrons. The exchange potential takes the form of exact exchange and its action on an arbitrary orbital $\psi_{p}$ is given by
@@
Correlation is absent, unlike DFT which has the exchange-correlation functional $V_{\textrm{XC}}$. To calculate the excited-state properties of a system in most linear-response methods, we need extra unoccupied (virtual) orbitals. Self-consistency is not required here because the unoccupied orbitals do not contribute to the density or exchange potential of the system.
@@
This full TDHF formulation in equation 3 is a non-Hermitian eigenvalue problem, with matrix size twice the number of transitions considered. For real orbitals, the equation can be transformed into a Hermitian problem of half the matrix size [24]. One widely used approximation (applicable for complex orbitals too) is neglecting the matrix $B$ in to equation 3 completely and only solving for $AX=\omega X$, which is again a Hermitian problem of half the matrix size. This Tamm-Dancoff approximation (TDA) not only makes the calculation less computationally intensive but also has been shown to perform better than full TDHF [25]. It is equivalent to Configuration Interaction Singles (CIS) [9]. While TDA has been known to underestimate the transition dipole moments [26] compared to the full TDHF, it gives good estimate of the excitation energies and molecular properties [27]. Neglecting the exchange and Hartree integrals in equation 6 and calculating the excitation energy as the difference of virtual and occupied eigenvalues is the IPA. This work will focus on TDA-TDHF, adapting Octopus’ existing implementation [19] of the transformed TDDFT equations for real orbitals (Casida equations) [24] and of Tamm-Dancoff TDDFT [28].
Once the TDHF eigenvalue problem is solved, one can obtain the absorption spectrum too. Using the TDA eigenvectors $\{x_{k,ia}\}$, the many-body dipole transition matrix elements reduce to a linear combination of single-particle matrix elements [22]:
@@
like in TDDFT, where the index $ia$ runs over the pairs of occupied-unoccupied orbitals. The isotropically averaged oscillator strength is then calculated as
@@ 3.1 Exchange Potential and Adaptively Compressed Exchange (ACE)
Unlike semi-local exchange and correlation density functionals such as the Local Density Approximation (LDA) [29, 30], the Hartree-Fock exchange potential is dependent on the molecular orbitals (equation 2) instead of the density. The non-local nature of this integral makes it time-consuming to solve numerically, especially in methods with large basis sets. Therefore real-space and plane-wave implementations use iterative algorithms which only require the application of $\hat{V}_{\rm X}$ to specific orbitals instead of the full construction of the operator $\hat{V}_{X}$.
Equation 2 can be solved as a Poisson equation with $\rho(r)=\psi_{j}(r)\psi_{i}(r)$, using various optimized Poisson solvers such as interpolating scaling functions, like for the Hartree potential [31]. The cost of solving equation 2 by direct integration for each $r$ is $\mathcal{O}(N_{\mu}^{2})$ while the cost of solving as a Poisson problem is $\mathcal{O}(N_{\mu}\log N_{\mu})$ where $N_{\mu}$ is the number of grid points. This is a significant improvement, but exact exchange still remains considerably more computationally intensive than a semi-local functional.
Adaptively Compressed Exchange (ACE) [32] is an approach we will use which was introduced to speed up the calculation of exact exchange. This method was previously implemented in Octopus. Other approaches available to speed up the exchange calculation include the Selected Column Density Matrix [33] and Recursive Subspace Bisection [21] methods. In the ACE formalism, the full-rank exchange operator, $\hat{V}_{\rm X}$, is approximated by the low rank ACE operator, $\hat{V}_{\rm X}^{\rm ACE}=\zeta\zeta^{T}$, which is obtained iteratively as follows.
@@
Once the solution to the ground-state is achieved, we calculate a set of unoccupied orbitals via a non-self-consistent calculation. ACE is used here too with the only difference from the ground-state calculation being the inclusion of unoccupied orbitals in equations 10, 11 and 12. Although one can obtain these unoccupied orbitals in the ground-state SCF by calculating an extra set of eigenfunctions with zero occupation which do not contribute to the total energy, it is preferable to calculate the ground-state without unoccupied orbitals first, since calculation of these extra eigenfunctions in all iterations before convergence is not useful. The number of unoccupied orbitals one can obtain also depends on the size of the basis set: the number of eigenfunctions obtainable is equal to the size of the Hamiltonian matrix. Hence, the number of unoccupied orbitals obtainable from GTOs will be lower than that obtainable from the (orders of magnitude) larger number of real-space grid points. Usually, one calculates all possible unoccupied orbitals with GTOs, but in real space only a small fraction of the possible unoccupied orbitals are calculated.
ACE calculates the ground-state exchange potential iteratively as a whole by reducing the number times the exchange matrix is calculated in the SCF cycle. It is not implemented yet in Octopus to calculate individual exchange integrals with ACE, as would be needed for its use in TDHF. Refs. [34, 35] have shown this is possible with plane-waves and Gaussian basis sets using discrete variable representation [36] schemes. Whether this provides any speedup in real-space is yet to be studied and is a subject offuture work. SCF, without ACE, requires solving $N_{e}^{2}$ integrals for each iteration. However TDA-TDHF requires solving $N_{e}N_{\rm unocc}$ integrals only once to create the TDHF matrix which is diagonalized. This means the prefactor for TDHF integrals is generally lower than that for the SCF problem, and so it is more necessary to reduce the cost of exchange in SCF than in TDHF.
@@ 3.2 Mixing in SCF cycle
The nonlocality of the exchange potential makes different and more difficult SCF convergence behavior than in DFT. To facilitate faster convergence, different quantities can be mixed together from consecutive iterations and be used in the next iteration. The quantity that is usually mixed is the density, but one can also mix the potential or the wavefunctions. It should also be noted that mixing non-local potential is not as straightforward as mixing local potentials. In our calculations, we mix only the local part of the potential, i.e. the Hartree potential $V_{\rm H}$.
@@ 3.3 Pseudopotentials
In real-space or plane-wave calculations, pseudopotentials (PPs) are the standard approach to account for the core electrons of a system only implicitly, which can be done because the core electrons usually do not take part in bonding or electronic and optical properties. In real-space, PPs also solve a crucial problem. The $-Z/r$ Coulomb potential is divergent at $r=0$ and makes the wavefunctions vary very rapidly near the nucleus, requiring an infeasible large number of grid points to describe. PPs circumvent this problem [17]. In Gaussian basis sets, usually all the electrons are treated explicitly, except for heavy elements in which case effective core potentials [40] are used, which are a form of PPs. PPs are generated from all-electron calculations for a given DFT functional. Typically only local or semi-local functionals are used. Approaches to generate PPs for HF have been developed [41, 42] but such PPs are rarely used in practice [43] and are not supported by Octopus. In this paper, we use optimized norm-conserving Vanderbilt [44] scalar relativistic pseudopotentials from pseudo-dojo [45]. We use LDA [29, 30] pseudopotentials for all HF ground-state calculations, as well as the LDA calculations reported later. For DFT calculations with BLYP [46, 47] for our TDDFT comparison, we have used PBE PPs as BLYP and PBE are both GGA functionals and pseudo-dojo currently doesn’t have BLYP PPs. For these TDDFT calculations, we note that Octopus currently only supports LDA kernels and so the LDA kernel [29, 30] is used for our TDDFT calculations based on a BLYP ground state.
Using PPs generated with one functional for a DFT calculation with a different functional is a common practice in DFT, and ubiquitous for hybrid functionals. A recent study shows that the inconsistent use of PPs does not make very significant differences [48] in DFT. In our study, although we are doing HF calculations, HF exchange is only applied to the valence electrons and not to the core electrons.
@@ 4.1 Speed-up of HF with ACE
To test the efficiency of the ACE formalism, we can look at the average time required to complete one iteration in the SCF cycle. Table 1 shows that the time per iteration for LDA calculations is $1-3\%$ of the time per iteration for HF calculations without ACE, i.e. the direct solution of the Poisson equations for exchange. With ACE, however, we have achieved a speed up of at least 7 times, with maximum speed up of 14 times for formamide. We have also confirmed numerically identical results between calculations with and without ACE. We also studied the scaling of time per SCF iteration as a function of number of grid-points for LDA, HF without ACE, and HF with ACE, as shown in Figure 1. We see similar scaling in each case: the scaling should be $\mathcal{O}\left(N_{\mu}\right)$ for LDA and $\mathcal{O}\left(N_{\mu}\log N_{\mu}\right)$ for HF, which look similar on this scale. However, the speed up from ACE is dramatic. This is a big improvement in making the ground-state HF calculations practical: without ACE they were two orders of magnitude more expensive than LDA, but with ACE are only one order of magnitude more expensive.
@@
Without ACE With ACE Speed up
@@
Table 1: Time per SCF iteration in different molecules. Times are calculated, on a node with two Intel 28 core Xeon Gold 6330 processors, by taking an average over the first 10 iterations for HF and an average over all iterations for LDA.
Figure 1: Scaling of time per SCF iteration for different approaches, showing similar scaling for each but an improved pre-factor for HF when using ACE. Times are calculated as an average for the first 20 SCF iterations for nitrosomethane molecule. Data obtained for fixed number of orbitals by varying grid spacing by $0.01$ from $0.20$ to $0.09$ on a simulation domain constructed from union of spheres of radii $5$ around each atom.

### v2

@@ Abstract
Time-dependent Hartree-Fock (TDHF) is one of the fundamental post-Hartree-Fock (HF) methods to describe excited states. In its Tamm-Dancoff form, equivalent to Configuration Interaction Singles, it is still widely used and particularly applicable to big molecules where more accurate methods may be unfeasibly expensive. However, it is rarely implemented in real space, mostly because of the expensive nature of the exact-exchange potential in real space. Compared to widely used Gaussian-type orbitals (GTO) basis sets, real space often offers easier implementation of equations and more systematic convergence of Rydberg states, as well as favorable scaling, effective domain parallelization, flexible boundary conditions, and ability to treat model systems. We implemented TDHF in the Octopus real-space code as a step toward linear-response hybrid time-dependent density-functional theory (TDDFT), other post-HF methods, and ensemble density-functional theory methods involving exact exchange. Calculation of HF’s non-local exact exchange is very expensive in real space. We overcome this limitation with Octopus’ implementation of Adaptively Compressed Exchange (ACE), and find the appropriate mixing scheme and starting point to complete the ground-state calculation in a practical amount of time, and thus enable TDHF. We compared our results to those from GTOs on a set of small molecules and confirmed close agreement of results, though with larger deviations than in the case of semi-local TDDFT. We find that convergence of TDHF demands a finer real-space grid than semi-local TDDFT. We also present the subtleties in benchmarking a real-space calculation against GTOs, relating to Rydberg and vacuum states.
@@ 1 Introduction
Hartree-Fock (HF) theory [1, 2] is a main pillar in quantum chemistry, serving as the starting point for calculating the electronic structure of molecules. HF calculations provide molecular geometry, electron distribution, and ground-state properties [3]. Once the ground state has been calculated, excited states can be calculated using the ground-state results. A first approximation one can make for the excitation energies is the eigenvalue differences of the unoccupied and occupied states, termed the Independent Particle Approximation (IPA) [4]. Configuration Interaction (CI), Equation-of-Motion Coupled Cluster (EOM-CC), and Time-Dependent Hartree-Fock (TDHF) are examples of post-Hartree-Fock methods [5] that can provide more accurate excited states. The exact exchange potential used in HF is orbital-dependent while it is density-dependent in traditional semi-local Kohn-Sham Density Functional Theory (DFT). Hybrid functionals which are a mixture of these two types of exchange potentials are widely used [6, 7] to improve accuracy, so TDHF is a building block for time-dependent density-functional theory (TDDFT) with hybrids [8].
@@
Hartree-Fock theory provides a reasonably accurate description of the ground state of molecular systems, especially for small-to-moderately sized molecules and systems where electron correlation effects are not dominant [8]. Gaussian-type orbitals (GTOs) are typically used as basis sets for Hartree-Fock [9]. Implementation of Hartree-Fock in real space is rare, though a finite-element-type implementation has been published [10]. Using GTOs conveniently allows integrals to be done analytically. However, the nature of orbitals one can obtain in an electronic structure calculations depends on the type and size of the basis set used. The basis set is typically optimized for the ground state which can make description of unoccupied orbitals difficult [11]. This issue is particularly acute for Rydberg states which are highly excited states that are far from the atoms and weakly bound [12], but transitions from valence to Rydberg states can be important in optical properties [13]. To include these states, additional diffuse functions are commonly added to the basis sets [14]. Comparisons of GTOs and real space have shown that such very large, diffuse basis sets can be needed to describe the excited states involved in non-linear optical response [15]. In real space, Rydberg states do not require special diffuse functions but can be captured by systematically converging the physical size of the real-space domain in which the orbitals are calculated. This domain size and the spacing of the grid-points are the parameters controlling the accuracy of a real-space calculation, which can be systematically converged [16]. Other advantages of real-space calculations include favorable scaling for large systems, effective parallelization over domains, and straightforward use of both finite and periodic boundary conditions [16, 17, 18], as well as the ability to handle non-atomic model systems [19]. In real-space however, the calculation of exchange integrals becomes very expensive [20] and these integrals are a major part of any HF calculation underlying TDHF.
@@
In this paper, we present results from our real-space implementation of TDHF in the Octopus code. In section 2, we discuss the general theory behind ground-state HF and TDHF. Section 3.1 discusses the exact exchange potential which is the main feature of HF which sets it apart from traditional Kohn-Sham DFT. As this is the most expensive part of the calculation, we discuss the Adatively Compressed Exchange (ACE) formalism to speed up the exact exchange calculation. We also present the choice of parameters used in our calculations in sections 3.3 and 3.4. In section 4 we discuss the general difficulties in reaching a solution to the HF ground-state in real-space and how to overcome them. Finally, section 4.5 shows the results of TDHF calculations, with comparison to excitation energies obtained from GTOs, demonstrating a good level of agreement and practical attainability of convergence with respect to the real-space grid parameters.
@@ 2 Theory
The four terms in the parentheses are the kinetic energy operator, external potential, Hartree potential, and exchange potential. The Hartree potential depends on the density $\rho\left(r\right)=\sum_{j}^{N_{e}}\left|\psi_{j}\right|^{2}$, where $N_{e}$ is the number of electrons. The exchange potential is used in its exact form and its action on an arbitrary orbital $\psi_{p}$ is given by
@@
Correlation is absent, unlike DFT which has the exchange-correlation functional $V_{\textrm{XC}}$. To calculate the excited-state properties of a system in most linear-response methods, we need extra unoccupied (virtual) orbitals. Self-consistency is not required in generating these unoccupied orbitals because they do not contribute to the density or exchange potential of the system.
@@
This full TDHF formulation in equation 3 is a non-Hermitian eigenvalue problem, with matrix size twice the number of transitions considered. For real orbitals, the equation can be transformed into a Hermitian problem of half the matrix size [23]. One widely used approximation (applicable for complex orbitals too) is neglecting the matrix $B$ in equation 3 completely and only solving $AX=\omega X$, which is again a Hermitian problem of half the matrix size. This Tamm-Dancoff approximation (TDA) applied to TDHF is equivalent to Configuration Interaction Singles (CIS) [8]. TDA not only makes the calculation less computationally intensive, but also can perform better than full TDHF [24] in many cases. Neglecting the exchange and Hartree integrals in equation 6 and calculating the excitation energy as the difference of virtual and occupied eigenvalues is the IPA. This work will focus on TDA-TDHF, adapting Octopus’ existing implementation [18] of HF and Tamm-Dancoff TDDFT [25].
Once the TDHF eigenvalue problem is solved, one can obtain the absorption spectrum too. Using the TDA singlet excitation eigenvectors $\{x_{k,ia}\}$, the many-body dipole transition matrix elements reduce to a linear combination of single-particle matrix elements [21]:
@@
like in TDDFT, where the index $ia$ runs over the pairs of occupied-unoccupied orbitals. The isotropically averaged oscillator strength, in the electric dipole approximation, is then calculated as
@@ 3.1 Exchange Potential and Adaptively Compressed Exchange (ACE)
Unlike semi-local exchange and correlation density functionals such as the Local Density Approximation (LDA) [26, 27], the Hartree-Fock exchange potential is dependent on the molecular orbitals (equation 2) instead of the density. The non-local nature of this integral makes it time-consuming to solve numerically, especially in methods with large basis sets. Therefore real-space and plane-wave implementations [28] use iterative algorithms which only require the application of $\hat{V}_{\rm X}$ to specific orbitals instead of the full construction of the operator $\hat{V}_{X}$.
Equation 2 can be solved as a Poisson equation with $\rho(r)=\psi_{j}(r)\psi_{i}(r)$, using various optimized Poisson solvers such as interpolating scaling functions, like for the Hartree potential [29]. The cost of solving equation 2 by direct integration for each $r$ is $\mathcal{O}(N_{\mu}^{2})$, while the cost of solving as a Poisson problem is $\mathcal{O}(N_{\mu}\log N_{\mu})$, where $N_{\mu}$ is the number of grid points. This is a significant improvement, but exact exchange still remains considerably more computationally intensive than a semi-local functional.
We will use the Adaptively Compressed Exchange (ACE) [30] approach, which was introduced to speed up the calculation of exact exchange. This method was previously implemented in Octopus. ACE is closely related to the density-fitting technique commonly used with GTOs in quantum chemistry [31]. Other possible approaches to speed up the exchange calculation include the Selected Column Density Matrix [32] and Recursive Subspace Bisection [20] methods. In the ACE formalism, the full-rank exchange operator $\hat{V}_{\rm X}$ is approximated by the low-rank ACE operator, $\hat{V}_{\rm X}^{\rm ACE}=\zeta\zeta^{\rm T}$, which is obtained iteratively as follows.
@@
Once the solution to the ground-state is achieved, we calculate a set of unoccupied orbitals via a non-self-consistent calculation. ACE is used here too, with the only difference from the ground-state calculation being the inclusion of unoccupied orbitals in equations 10, 11 and 12. Although one can obtain these unoccupied orbitals in the ground-state SCF by calculating an extra set of eigenfunctions with zero occupation which do not contribute to the total energy, it is preferable to calculate the ground-state without unoccupied orbitals first, since calculation of these extra eigenfunctions in all iterations before convergence is not useful. The number of unoccupied orbitals one can obtain also depends on the size of the basis set: the number of eigenfunctions obtainable is equal to the size of the Hamiltonian matrix. Hence, the number of unoccupied orbitals obtainable from GTOs will be lower than that obtainable from the (orders of magnitude) larger number of real-space grid points. Usually, one calculates all possible unoccupied orbitals with GTOs, but in real space only a small fraction of the possible unoccupied orbitals are calculated.
ACE calculates the ground-state exchange potential iteratively as a whole by reducing the number times the exchange matrix is calculated in the SCF cycle. It is not implemented yet in Octopus to calculate individual exchange integrals with ACE, as would be needed for its use in TDHF. Refs. [33, 34] have shown this is possible with plane-waves and Gaussian basis sets using discrete variable representation [35] schemes. Whether this provides any speedup in real-space is yet to be studied and is a subject for future work. SCF, without ACE, requires solving $N_{e}^{2}$ integrals for each iteration. However, TDA-TDHF requires solving $N_{e}N_{\rm unocc}$ integrals only once, to create the TDHF matrix which is diagonalized. This means the prefactor for TDHF integrals is generally lower than that for the SCF problem, and so it is more necessary to reduce the cost of exchange in SCF than in TDHF.
@@ 3.2 Mixing in SCF cycle
The nonlocality of the exchange potential makes different and more difficult SCF convergence behavior than in DFT. To facilitate faster convergence, different quantities can be mixed together from consecutive iterations and be used in the next iteration. The quantity that is usually mixed is the density, but one can also mix the potential or the wavefunctions. It should also be noted that mixing non-local potentials is not as straightforward as mixing local potentials. In our calculations with potential mixing, we mix only the local part of the potential, i.e. the Hartree potential $V_{\rm H}$.
@@ 3.3 Pseudopotentials
In real-space or plane-wave calculations, pseudopotentials (PPs) are the standard approach to account only implicitly for the core electrons of a system, which can be done because the core electrons usually do not take part in bonding or electronic and optical properties. In real-space, PPs also solve a crucial problem. The $-Z/r$ Coulomb potential is divergent at $r=0$ and makes the wavefunctions vary very rapidly near the nucleus, requiring an infeasible large number of grid points to describe. PPs circumvent this problem [16]. By contrast, in Gaussian basis sets, usually all the electrons are treated explicitly, except for heavy elements in which case effective core potentials (a form of PPs) [39] are used. PPs are generated from all-electron calculations for a given DFT functional. Typically only local or semi-local functionals are used. Approaches to generate PPs for HF have been developed [40, 41] but such PPs are rarely used in practice [42] and are not supported by Octopus. In this paper, we use optimized norm-conserving Vanderbilt [43] scalar relativistic pseudopotentials from pseudo-dojo [44]. We use LDA [26, 27] pseudopotentials for all HF ground-state calculations, as well as the LDA calculations reported later. For DFT calculations with BLYP [45, 46] for our TDDFT comparison, we have used PBE PPs as BLYP and PBE are both GGA functionals and pseudo-dojo currently doesn’t have BLYP PPs. For these TDDFT calculations, we note that Octopus currently only supports LDA kernels and so the LDA kernel [26, 27] is used for our TDDFT calculations based on a BLYP ground state, as is often done in Octopus calculations [15]. Using PPs generated with one functional for a DFT calculation with a different functional is a common practice in DFT, and ubiquitous for hybrid functionals. A recent study shows that the inconsistent use of PPs does not make very significant differences [47] in DFT.
@@ 4.1 Speed-up of HF with ACE
To test the efficiency of the ACE formalism, we can look at the average time required to complete one iteration in the SCF cycle. Table 1 shows that the time per iteration for LDA calculations is $1-3\%$ of the time per iteration for HF calculations without ACE, i.e. the direct solution of the Poisson equations for exchange. With ACE, however, we have achieved a speed-up of at least 7 times, with maximum speed-up of 14 times for formamide. We have also confirmed numerically identical results between calculations with and without ACE. We also studied the scaling of time per SCF iteration as a function of number of grid-points for LDA, HF without ACE, and HF with ACE, as shown in Figure 1. We see similar scaling in each case: the scaling should be $\mathcal{O}\left(N_{\mu}\right)$ for LDA and $\mathcal{O}\left(N_{\mu}\log N_{\mu}\right)$ for HF, which look similar on this scale. However, the speed-up due to ACE is dramatic. This is a big improvement in making the ground-state HF calculations practical: without ACE they were two orders of magnitude more expensive than LDA, but with ACE are only one order of magnitude more expensive.
@@
Without ACE With ACE Speed-up
@@
Table 1: Time per SCF iteration in different molecules, for runs in parallel using a node with two Intel 28 core Xeon Gold 6330 processors. Times are calculated by taking an average over the first 10 iterations for HF and an average over all iterations for LDA.
Figure 1: Scaling of time per SCF iteration for different approaches, showing similar scaling for each but an improved pre-factor for HF when using ACE. Times are calculated as an average for the first 20 SCF iterations for the nitrosomethane molecule. Data was obtained for a fixed number of orbitals by varying the grid spacing by $0.01$ from $0.20$ to $0.09$ on a simulation domain constructed from a union of spheres of radii $5$ around each atom.